# The Nature of Optical Afterglows Without Gamma-ray Bursts: Identification of AT2023lcr and Multiwavelength Modeling

### v1

@@ 1 Introduction
In this paper, we present the identification, follow-up, and multi-wavelength modeling of one of the most recent such events, AT 2023lcr. We only consider the afterglow light curve, although an associated Ic-BL supernova was identified at a later time (Martin-Carrillo et al., 2023a). As shown in Table 1, AT 2023lcr is one of six afterglows discovered in optical-survey data with no detected GRB but with a measured redshift. To put our AT 2023lcr results into context, we also present multi-wavelength modeling of three afterglows in Table 1: AT 2020blt, AT 2021any, and AT 2021lfa. For AT 2019pim, we refer the reader to Perley et al. (2024), who used a similar approach to this work, while AT 2023sva will be modeled in future work (Srinivasaragavan et al. in prep). All afterglows are modeled using Markov Chain Monte Carlo (MCMC) methods in emcee (version 3.1.4; Foreman-Mackey et al. 2013) and afterglow models from afterglowpy (version 0.7.3; Ryan et al. 2020).
@@
AT 2023sva 2.281 [18, 19, 20, 21] -
References. — [0] this work, [1] Cenko et al. (2013), [2] Wang & Dai (2013), [3] Wu et al. (2013), [4] Perley et al. (2024), [5] Ho et al. (2020), [6] Sarin et al. (2022), [7] Ho et al. (2022), [8] Gupta et al. (2022), [9] Xu et al. (2023), [10] Lipunov et al. (2022), [11] Ye et al. (2024), [12] Wang et al. (2023), [13] Ho (2023), [14] Andreoni et al. (2023), [15] Perley (2023), [16] Vail et al. (2023a), [17] Sfaradi et al. (2023), [18] Vail et al. (2023b), [19] de Ugarte Postigo et al. (2023), [20] Rhodes et al. (2023), [21] Roberts et al. (2023).
@@
AT 2020blt was previously modeled in Sarin et al. (2022) (hereafter S22); AT 2021any was previously modeled in Gupta et al. (2022) (G22) and Xu et al. (2023) (X23); AT 2021lfa was previously modeled in Ye et al. (2024) (Y24). In this work, we explore additional jet structures, modeling configurations, and constraints on afterglow behavior, discussing comparisons between the mentioned works and this work in Section 5.
@@ 2.1 Optical Photometry
AT 2023lcr was initially reported (Tonry et al., 2023; Fulton et al., 2023) to the Transient Name Server (TNS) by the ATLAS survey (Tonry et al., 2018) as ATLAS23msn. AT 2023lcr was also detected as part of the ZTF high-cadence partnership survey (Bellm et al., 2019b), with the first detection at [^1] All times in this paper are in UT. 06:36:27 on 2023 June 18 (internal name ZTF23aaoohpy) at a position $\alpha=16$:31:37.416 and $\delta=+$26:21:58.31 (J2000) and a Galactic latitude $b=41.25$ deg (Swain et al., 2023a). AT 2023lcr was flagged as a transient of interest due to its rapid rise ($>1.2\,$mag d-1 in $r$ band [^2] All magnitudes are in AB unless specified otherwise.), red colors ($g-r=0.29\pm 0.08\,$mag, corrected for Milky Way extinction assuming $R_{v}=3.1$; Schlegel et al. 1998), and lack of a bright host-galaxy counterpart.
@@ 2.3 X-rays
AT 2023lcr was also observed by the Advanced CCD imaging spectrometer (ACIS; Garmire et al. 2003) on board the Chandra X-Ray Observatory (Chandra) [^8] The data is contained in https://doi.org/10.25574/27933 (catalog ObsID 27933). under a Director’s Discretionary Time proposal [^9] Proposal Number 24508916, PI A. Martin-Carrillo., four days after the final Swift/XRT detection. We reduced the data using the Chandra Interactive Analysis of Observations (CIAO; Fruscione et al. 2006) software package (v4.15). Counts were extracted from AT 2023lcr using a circle with radius $2^{\prime\prime}$, and background counts were measured in source-free regions near AT 2023lcr. We used specextract to bin the spectrum (with 5 counts per bin). The routine sherpa was used to fit the spectrum in the range 0.5–6 keV, with the background subtracted, using a model with photoelectric absorption and a single-component power law (xsphabs.abs1 $\times$ powlaw1d.p1). We set the Galactic hydrogen column density to be the same as for the Swift observations. The best-fit power law index was $\Gamma=1.16^{+0.95}_{-0.95}$ (68% confidence), consistent with the value from Swift but with much larger uncertainties, so for consistency we also adopt $\Gamma=2$. The flux reported in Table 2 has been multiplied by a factor of 1.77 to correct from the 0.5–6 keV band to the 0.3–10 keV band. In Table 2, we also present the spectral flux density at 1 keV assuming a spectral index of $f_{\nu}\propto\nu^{-1}$.
@@ 3.1 Optical
where $m(t)$ is the apparent magnitude as a function of time, $n$ parameterizes the smoothness of the break ($n=\infty$ is a sharp break), $\alpha_{1}$ and $\alpha_{2}$ are the pre- and post-break temporal indices respectively, $t_{0}$ is the time of the explosion in days, $t_{b}$ is the time of the break in days, and $m_{c}$ is the magnitude at the time of the break assuming $n=\infty$. We assume no contribution from the host galaxy or the underlying SN, because we do not observe any flattening in the optical light curve; the accompanying Ic-BL SN was likely too faint to make a significant contribution at the time of the presented optical observations. We fixed $n=2$ and allowed $t_{0}$ to vary between the time of the last GOTO non-detection and the first GOTO detection.
@@ 3.2 X-rays
We fit a single power law to the X-ray light curve of AT 2023lcr at 1 keV, and find a best-fit power law index $\alpha_{X}=1.47^{+0.17}_{-0.16}$ (68% confidence) where $f_{\nu}\propto t^{-\alpha_{X}}$, consistent with the slope of the optical light curve in the same time range. We assume no host galaxy extinction. There was only one X-ray detection of AT 2021lfa, at a similar flux to AT 2023lcr. The X-ray light curves of AT 2023lcr and AT 2019pim have similar temporal decay indices, but the X-ray flux density of AT 2019pim was an order of magnitude fainter.
@@ 4.1 Settings
We use afterglowpy (version 0.7.3; Ryan et al. 2020) and MCMC methods in emcee (version 3.1.4; Foreman-Mackey et al. 2013) to fit the radio, optical, and X-ray observations of AT 2023lcr with a set of physical parameters that describe the jet and circumburst medium. To put AT 2023lcr’s modeling into context, we also perform a consistent analysis on three similarly discovered events, shown in Table 1: AT 2020blt, AT 2021any, and AT 2021lfa. For AT 2019pim, see Perley et al. (2024), which explored a similarly broad range of scenarios; AT 2023sva will be modeled in future work (Srinivasaragavan et al, in prep.). By contrast, previous works modeling AT 2020blt (Sarin et al., 2022), AT 2021any (Gupta et al., 2022; Xu et al., 2023), and AT 2021lfa (Ye et al., 2024) explored more fixed setups.
All data used in this work for AT 2020blt and AT 2021any can be found in Ho et al. (2020, 2022). To model AT 2021lfa, we used all observations from Ho et al. (2022) and additional observations from Lipunov et al. (2022). Optical observations were corrected for Galactic extinction assuming diffuse reddening $R_{v}=3.1$ (Schlegel et al., 1998). For all afterglows, we converted X-ray fluxes to a 5 keV flux density assuming a spectral index of $\beta=-1$, where $f_{\nu}\propto\nu^{\beta}$.
@@ 5.2 AT2023lcr
Figure 5: On-axis top hat jet with $\Gamma_{0}\approx 166$ for AT 2023lcr. Plotted are light curves generated from 150 randomly selected posterior samples. Radio upper limits are plotted at 3 $\times$ image RMS. The samples are generally consistent with observations, but overestimate radio L-band (1.39 GHz) and S-band (3.5, 3.75 GHz) detections, likely due to afterglowpy’s lack of self-absorption modeling.
@@ 5.3 AT2020blt
Figure 6: On-axis top hat jet with $\Gamma=\infty$ for AT 2020blt. Plotted are light curves generated from 150 randomly selected posterior samples. Radio upper limits are plotted at 3 $\times$ image RMS. Samples are consistent with optical and radio observations, but overestimate the X-ray detection by $\sim 1.5$ orders of magnitude, possibly due to unmodeled central engine energy injection.
@@
Figure 7: The lowest-likelihood off- and on-axis samples for AT 2020blt. As expected, the off-axis model has a later peak; otherwise, the samples produce similar fits. Radio upper limits are plotted at 3 $\times$ image RMS.
@@ 5.4 AT2021any
Figure 8: Left: On-axis top hat jet with $\Gamma_{0}\approx 204$ for AT 2021any where the burst time was allowed to vary. Right: On-axis top hat jet with $\Gamma_{0}\approx 81$ and a fixed $t_{0}=59230.290$ MJD. Plotted are light curves generated from 150 randomly selected posterior samples. Radio upper limits are plotted at 3 $\times$ image RMS. Both models are consistent with optical and X-ray observations, but struggle with radio X-band detections, possibly due to interstellar scintillation.
@@
Figure 9: Top: On-axis top hat jet with $\Gamma_{0}\approx 11.5$ for AT 2021lfa. The model is consistent with optical and radio observations, but underestimates the X-ray detection by an order of magnitude. Bottom: On-axis Gaussian jet with $\Gamma_{0}\approx 6.0$ for AT 2021lfa, which is consistent with all observations. Plotted are light curves generated from 150 randomly selected posterior samples.
@@ 5.5.2 Off-axis Interpretation
Figure 10: Top: Off-axis Gaussian jet with $\Gamma_{0}\approx 110$ for AT 2021lfa, which struggles with the finer features of the rising phase detections. Bottom: Off-axis power law jet with $\Gamma_{0}\approx 155$ for AT 2021lfa, which is consistent with all optical observations. Both the Gaussian and power law models overestimate the X-ray detection and late-time radio observations. Plotted are light curves generated from 150 randomly selected posterior samples.
@@ Appendix A Optical data
Start Date (MJD) Instrument Filter Maga
@@
aafootnotetext: Not corrected for Milky Way extinction
@@ Appendix B Radio data
Epoch Start MJD Instrument $\nu_{\mathrm{obs}}$ (GHz) $f_{\nu}$ ($\mu$Jy)
@@
aafootnotetext: Upper limits reported as 3$\times$ image RMS. Epochs of VLA observations are numbered. We report statistical errors, but include systematic errors in our reported reduced $\chi^{2}$/DoFs.

### v2

@@ 1 Introduction
In this paper, we present the identification, follow-up, and multi-wavelength modeling of one of the most recent such events, AT 2023lcr. We only consider the afterglow light curve, although an associated Ic-BL supernova was identified at a later time (Martin-Carrillo et al., 2023a). As shown in Table 1, AT 2023lcr is one of six afterglows discovered in optical-survey data with no detected GRB but with a measured redshift. To put our AT 2023lcr results into context, we also present multi-wavelength modeling of three afterglows in Table 1: AT 2020blt, AT 2021any, and AT 2021lfa. For AT 2019pim, we refer the reader to Perley et al. (2024), who used a similar approach to this work; for AT 2023sva, we refer the reader to (Srinivasaragavan et al., 2025). All afterglows are modeled using Markov Chain Monte Carlo (MCMC) methods in emcee (version 3.1.4; Foreman-Mackey et al. 2013) and afterglow models from afterglowpy (version 0.7.3; Ryan et al. 2020).
@@
AT 2023sva 2.281 [18, 19, 20, 21, 22] slightly off-axis structured jet [22]
References. — [0] this work, [1] Cenko et al. (2013), [2] Wang & Dai (2013), [3] Wu et al. (2013), [4] Perley et al. (2024), [5] Ho et al. (2020), [6] Sarin et al. (2022), [7] Ho et al. (2022), [8] Gupta et al. (2022), [9] Xu et al. (2023), [10] Lipunov et al. (2022), [11] Ye et al. (2024), [12] Wang et al. (2023), [13] Ho (2023), [14] Andreoni et al. (2023), [15] Perley (2023), [16] Vail et al. (2023a), [17] Sfaradi et al. (2023), [18] Vail et al. (2023b), [19] de Ugarte Postigo et al. (2023), [20] Rhodes et al. (2023), [21] Roberts et al. (2023), [22] Srinivasaragavan et al. (2025).
@@
AT 2020blt was previously modeled in Sarin et al. (2022) (hereafter S22); AT 2021any was previously modeled in Gupta et al. (2022) (G22) and Xu et al. (2023) (X23); AT 2021lfa was previously modeled in Ye et al. (2024) (Y24). In this work [^1] Code and data products can be found in the repository https://github.com/liluhua2/afterglowfit-public., we explore additional jet structures, modeling configurations, and constraints on afterglow behavior, discussing comparisons between the mentioned works and this work in Section 5.
@@ 2.1 Optical Photometry
AT 2023lcr was initially reported (Tonry et al., 2023; Fulton et al., 2023) to the Transient Name Server (TNS) by the ATLAS survey (Tonry et al., 2018) as ATLAS23msn. AT 2023lcr was also detected as part of the ZTF high-cadence partnership survey (Bellm et al., 2019b), with the first detection at [^2] All times in this paper are in UT. 06:36:27 on 2023 June 18 (internal name ZTF23aaoohpy) at a position $\alpha=16$:31:37.416 and $\delta=+$26:21:58.31 (J2000) and a Galactic latitude $b=41.25$ deg (Swain et al., 2023a). AT 2023lcr was flagged as a transient of interest due to its rapid rise ($>1.2\,$mag d-1 in $r$ band [^3] All magnitudes are in AB unless specified otherwise.), red colors ($g-r=0.29\pm 0.08\,$mag, corrected for Milky Way extinction with $E(B-V)=A_{V}/R_{V}=0.05$, assuming $R_{v}=3.1$; Schlegel et al. 1998), and lack of a bright host-galaxy counterpart.
@@ 2.3 X-rays
AT 2023lcr was also observed by the Advanced CCD imaging spectrometer (ACIS; Garmire et al. 2003) on board the Chandra X-Ray Observatory (Chandra) [^9] This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in Chandra Data Collection https://doi.org/10.25574/cdc.364 (catalog doi:10.25574/cdc.364). under a Director’s Discretionary Time proposal [^10] Proposal Number 24508916, PI A. Martin-Carrillo., four days after the final Swift/XRT detection. We reduced the data using the Chandra Interactive Analysis of Observations (CIAO; Fruscione et al. 2006) software package (v4.15). Counts were extracted from AT 2023lcr using a circle with radius $2^{\prime\prime}$, and background counts were measured in source-free regions near AT 2023lcr. We used specextract to bin the spectrum (with 5 counts per bin). The routine sherpa (Freeman et al., 2001; Doe et al., 2007) was used to fit the spectrum in the range 0.5–6 keV, with the background subtracted, using a model with photoelectric absorption and a single-component power law (xsphabs.abs1 $\times$ powlaw1d.p1). We set the Galactic hydrogen column density to be the same as for the Swift observations. The best-fit power law index was $\Gamma=1.16^{+0.95}_{-0.95}$ (68% confidence), consistent with the value from Swift but with much larger uncertainties, so for consistency we also adopt $\Gamma=2$. The flux reported in Table 2 has been multiplied by a factor of 1.77 to correct from the 0.5–6 keV band to the 0.3–10 keV band. In Table 2, we also present the spectral flux density at 1 keV assuming a spectral index of $f_{\nu}\propto\nu^{-1}$.
@@ 3.1 Optical
where $m(t)$ is the apparent magnitude as a function of time, $n$ parameterizes the smoothness of the break ($n=\infty$ is a sharp break), $\alpha_{1}$ and $\alpha_{2}$ are the pre- and post-break temporal indices respectively, $t_{0}$ is the time of the explosion in days, $t_{b}$ is the time of the break in days, and $m_{c}$ is the magnitude at the time of the break assuming $n=\infty$. We assume no contribution from the host galaxy or the underlying SN. A typical GRB-SN would be both too faint for our observations and redshifted to the NIR. These assumptions are supported by the lack of observed flattening in the optical light curve. We fixed $n=2$ and allowed $t_{0}$ to vary between the time of the last GOTO non-detection and the first GOTO detection.
@@ 3.2 X-rays
We fit a single power law to the X-ray light curve of AT 2023lcr at 1 keV, and find a best-fit power law index $\alpha_{X}=1.47^{+0.17}_{-0.16}$ (68% confidence) where $f_{\nu}\propto t^{-\alpha_{X}}$, consistent with the slope of the optical light curve in the same time range. Given the similar optical and X-ray spectral indices ($f_{\nu}\propto\nu^{-1}$), we assume no host galaxy extinction. There was only one X-ray detection of AT 2021lfa, at a similar flux to AT 2023lcr. The X-ray light curves of AT 2023lcr and AT 2019pim have similar temporal decay indices, but the X-ray flux density of AT 2019pim was an order of magnitude fainter.
@@ 4.1 Settings
We use afterglowpy (version 0.7.3; Ryan et al. 2020) and MCMC methods in emcee (version 3.1.4; Foreman-Mackey et al. 2013) to fit the radio, optical, and X-ray observations of AT 2023lcr with a set of physical parameters that describe the jet and circumburst medium. To put AT 2023lcr’s modeling into context, we also perform a consistent analysis on three similarly discovered events, shown in Table 1: AT 2020blt, AT 2021any, and AT 2021lfa. For AT 2019pim, see Perley et al. (2024), which explored a similarly broad range of scenarios; for AT 2023sva, see (Srinivasaragavan et al., 2025). By contrast, previous works modeling AT 2020blt (Sarin et al., 2022), AT 2021any (Gupta et al., 2022; Xu et al., 2023), and AT 2021lfa (Ye et al., 2024) explored more fixed setups.
All data used in this work for AT 2020blt and AT 2021any can be found in Ho et al. (2020, 2022). To model AT 2021lfa, we used all observations from Ho et al. (2022) and additional observations from Lipunov et al. (2022). Optical observations were corrected for Galactic extinction with $E(B-V)=A_{V}/R_{V}$, assuming $R_{V}=3.1$ (Schlegel et al., 1998). For all afterglows, we converted X-ray fluxes to a 5 keV flux density assuming a spectral index of $\beta=-1$, where $f_{\nu}\propto\nu^{\beta}$.
@@ 5.2 AT2023lcr
Figure 5: On-axis top hat jet with $\Gamma_{0}\approx 166$ for AT 2023lcr, fit to X-ray, optical (left), and radio observations (right). The model is able to reproduce overall trends in all bands, especially the early-time optical $L$-band detection, the optical light curve break, and the X-ray observations. However, the model overestimates the radio L-band (1.39 GHz) and S-band (3.5, 3.75 GHz) detections, likely due to afterglowpy’s lack of self-absorption modeling. Plotted are light curves generated from 150 randomly selected posterior samples. Radio upper limits are plotted at 3 $\times$ image RMS.
@@ 5.3 AT2020blt
Figure 6: On-axis top hat jet with $\Gamma=\infty$ for AT 2020blt, fit to optical (left), X-ray, and radio observations (right). Plotted are light curves generated from 150 randomly selected posterior samples. The model is consistent with optical and radio observations, but underestimates the X-ray detection by $\sim 1.5$ orders of magnitude, possibly due to unmodeled central engine energy injection. Radio upper limits are plotted at 3 $\times$ image RMS.
@@
Figure 7: The lowest-likelihood off- and on-axis samples for AT 2020blt, fit to optical (top), X-ray, and radio observations (bottom). As expected, the off-axis model has a later peak; otherwise, the samples produce similar fits. Radio upper limits are plotted at 3 $\times$ image RMS.
@@ 5.4 AT2021any
Figure 8: Left: On-axis top hat jet with $\Gamma_{0}\approx 204$ for AT 2021any where the burst time was allowed to vary. Right: On-axis top hat jet with $\Gamma_{0}\approx 81$ and a fixed $t_{0}=59230.290$ MJD. Both models are consistent with optical and X-ray observations, but struggle with radio X-band detections, possibly due to interstellar scintillation. Plotted are light curves generated from 150 randomly selected posterior samples. Radio upper limits are plotted at 3 $\times$ image RMS.
@@
Figure 9: Top: On-axis top hat jet with $\Gamma_{0}\approx 11.5$ for AT 2021lfa, fit to X-ray, optical (left), and radio observations (right). The model is consistent with optical and radio observations, but underestimates the X-ray detection by an order of magnitude. Bottom: On-axis Gaussian jet with $\Gamma_{0}\approx 6.0$ for AT 2021lfa, which is consistent with all observations. Plotted are light curves generated from 150 randomly selected posterior samples.
@@ 5.5.2 Off-axis Interpretation
Figure 10: Top: Off-axis Gaussian jet with $\Gamma_{0}\approx 110$ for AT 2021lfa, fit to X-ray, optical (left), and radio observations (right). The model struggles with the finer features of the rising phase detections. Bottom: Off-axis power law jet with $\Gamma_{0}\approx 155$ for AT 2021lfa, which is consistent with all optical observations. Both the Gaussian and power law models overestimate the X-ray detection and late-time radio observations. Plotted are light curves generated from 150 randomly selected posterior samples.
@@ Appendix A Optical data
Start Date (MJD) Instrumenta Filter Magb
@@
aafootnotetext: P48: Palomar Observatory 48-inch Samuel Oschin Telescope; GIT: GROWTH-India Telescope; LT: Liverpool Telescope.
bbfootnotetext: Not corrected for Milky Way extinction
@@ Appendix B Radio data
Epoch Start MJD Instrumentb $\nu_{\mathrm{obs}}$ (GHz) $f_{\nu}$ ($\mu$Jy)
@@
aafootnotetext: Upper limits reported as 3$\times$ image RMS. Epochs of VLA observations are numbered. We report statistical errors, but include systematic errors in our reported reduced $\chi^{2}$/DoFs. bbfootnotetext: VLA: Karl G. Jansky Very Large Array; SMA: Submillimeter Array; ALMA: Atacama Large Millimeter/sub-millimeter Array; GMRT: Giant Metrewave Radio Telescope.